# Spinal open books and symplectic fillings with exotic fibers

### v1

@@ Abstract.
Spinal open book decompositions provide a natural generalization of open book decompositions. We show that any minimal symplectic filling of a contact $3$-manifold supported by a planar spinal open book is deformation equivalent to the complement of a positive multisection in a bordered Lefschetz fibration, which generalizes a result of Wendl [Wen10b]. Along the way, we give an explicit local model for a non-compactly supported singularity in a generalized version of bordered Lefschetz fibrations, given by pseudoholomorphic foliations associated to the spinal open books. This provides new tools to classify symplectic fillings of a contact $3$-manifold that is not supported by an amenable spinal open book, by studying monodromy factorizations in the newly defined spinal mapping class group. As an application, we complete the classification of strong fillings of all parabolic torus bundles, and make progress towards classifying symplectic fillings of contact $3$-manifolds supported by non-planar open books.
@@ 1. Introduction
Pseudoholomorphic foliations given by planar open book decompositions have been a powerful tool in studying classification problems of symplectic and Stein fillings of contact $3$-manifolds up to deformation. For example, see [Wen10b, PVHM10, Kal15, Pla12, Wan12, KL16] for applications. This paper further develops the tools of planar spinal open books and the pseudoholomorphic foliations induced by it and gives several applications to show their potential into solving classification problems previously inaccessible to simply using planar open books. Our main theorems are in §1.3. The study of spinal open book was initiated by Lisi, Van-Horn-Morris and Wendl in [LHMW18] and [LHMW20]. We discuss the current status and definition of spinal open books in §1.2 and §2.1.
@@ 1.1. Scientific context
The main results of [Wen10b], [LHMW18] and [LHMW20] encapsulate a general theme of classifying minimal symplectic fillings of contact $3$-manifolds. The primary idea is to find planar pages in the contact $3$-manifold, which are genus-zero pages of its open book decomposition and which can be lifted to pseudoholomorphic curves in the symplectization. One then considers the moduli space of these pseudoholomorphic curves which turn out to foliate the symplectic fillings of the contact $3$-manifold, as in the main theorems of [Wen10b, LHMW18, LHMW20]. In [LHMW18] and [LHMW20], which generalize the results of [Wen10b], the authors show that in general, the moduli spaces contain two types of degenerations, namely singular fibers and exotic fibers. The exotic fibers do not appear in the setting of planar open books of [Wen10b], and the authors of [LHMW18] and [LHMW20] also only classify symplectic fillings in the restricted Lefschetz-amenable setting, where exotic fibers do not appear. They show that, in this situation, all minimal strong fillings are deformation equivalent to Stein fillings. Furthermore, the pseudoholomorphic foliation gives a bordered Lefschetz fibration structure on the strong filling, which induce the planar spinal open book on the boundary. These results already give vast generalization of the relative case of a result of McDuff [McD90] and unify the results of Gromov [Gro85], Eliashberg [Eli91], and McDuff [McD91], which classified the minimal symplectic fillings of the standard $3$-sphere and the standard lens spaces. In addition, the main result of [Wen10b] turns the classification problem for $3$-manifolds supported by planar open books into a monodromy factorization problem of the mapping classes supported in the interior of the page. This reduction to monodromy factorization was used extensively in [PVHM10], [Kal15], [Pla12], [Wan12], and [KL16] to classify fillings of certain families of $3$-manifolds.
In this paper we extend the results above to turn the problem of classifying symplectic fillings of a general planar spinal open book into a monodromy factorization problem, in a larger mapping class group that we call the spinal mapping class group of a surface. This mapping class group in particular contains diffeomorphisms of a surface that are not supported in the interior and will be defined in §2.5. We show that the presence of exotic fibers in minimal symplectic fillings forces the fillings to be complements of multisections in bordered Lefschetz vibrations.
@@ 1.2. Past structural theorem for planar spinal open books
Though the nomenclature and formal definition appeared later via the collaborations [LHMW18, LHMW20], the notion of spinal open books was first studied by Wendl in [Wen10b], where he used them to classify the symplectic and Stein fillings of the standard $3$-torus.
@@
$\mathcal{M}=\mathcal{M}_{\operatorname{\text{reg}}}(\widehat{J})\cup\mathcal{M}_{\operatorname{\text{sing}}}(\widehat{J})\cup\mathcal{M}_{\operatorname{\text{exot}}}(\widehat{J}).$
@@
  • $u\in\Pi^{-1}(\mathcal{M}_{\operatorname{\text{reg}}}(\widehat{J}))$ is an embedded $\widehat{J}$-holomorphic curve asymptotic to simply covered Reeb orbits;
  • $u\in\Pi^{-1}(\mathcal{M}_{\operatorname{\text{sing}}}(\widehat{J}))$ is a nodal $\widehat{J}$-holomorphic curve asymptotic to simply covered Reeb orbits composing of two embedded curves intersecting transversely exactly once;
  • $u\in\Pi^{-1}(\mathcal{M}_{\operatorname{\text{exot}}}(\widehat{J}))$ is an embedded $\widehat{J}$-holomorphic curve with one asymptotic end doubly covering a Reeb orbit and all other ends simply covering Reeb orbits.
@@
The elements of the above moduli spaces are referred to as the moduli space of regular, singular and exotic fibers, respectively. The pseudoholomorphic foliations given by Theorem 1.1 in the symplectic filling naturally gives constraint to its boundary spinal open book. This leads to the seemingly technical definition of uniformity of a spinal open book, given in §2.1. As discussed in [LHMW20, Lemma 6.33], $\Pi|_{S_{i}}$ is an honest covering map if and only if $\mathcal{M}_{\operatorname{\text{exot}}}=\varnothing$. This condition on the spinal open book is called Lefschetz-amenable. In the Lefschetz-amenable setting, Lisi–Van-Horn-Morris–Wendl obtained the following result:
@@ 1.3. Main results
We extend the results in the previous section to strong fillings of contact structures supported by uniform, non-Lefschetz-amenable spinal open books. The main contribution of our paper is to study the mysterious exotic fibers naturally arising in the non-Lefschetz-amenable situation, give local models and use them to classify fillings. As a result, we obtain a strong fillability criterion for non-Lefschetz-amenable planar spinal open books. This enlarges the types of singularities in the literature that is usually considered in the classification problem for compact symplectic manifolds in a significant way, since these exotic singularities are in some sense a singularity “at infinity” and had not been extensively studied previously. See Remark 1.5 for other places where this type of singularity has been observed and relation to other works. In light of understanding these exotic fibers, we prove the following main result.
@@
We prove several technical results in §4 and §5 that allow us to conclude Theorem B. Firstly, Theorem 1.1 tells us that the completion of any strong filling of a partially planar spinal open book, consists of a partition by regular, singular and exotic fibers. We establish a local model for exotic fibers in Proposition 4.9, which gives us a count of exotic fibers in the filling, as well as the monodromy around an exotic fiber. In terms of spinal open books, we state the following theorem.
@@ 1.4. Symplectic fillings of torus bundles
Ever since Eliashberg [Eli96] determined the strong fillability of contact structures on $T^{3}$, there has been extensive studies on the symplectic fillability of torus bundles. In particular, Stein and strong fillability of contact structures on torus bundles are completely determined except for negative hyperbolic ones, through extensive studies by [DG01, EH02, Gay06, VHM07, GL16, DL18, Chr21]. In addition, several classification results for these torus bundles are also known, see [EH02, Chr21, Kal15, Lis08, LHMW20, PVHM10] for examples.
@@
where $\phi_{n}\colon\mathbb{R}/\mathbb{Z}\to\mathbb{R}$ satisfies
@@
±$\pm A_{k}=\pm\begin{pmatrix}1&k\\ 0&1\end{pmatrix}$
for some $k\in\mathbb{Z}$. We say a parabolic torus bundle $T_{A}$ is positive if $A$ is conjugate to $A_{k}$ and negative if $A$ is conjugate to $-A_{k}$. Let $T_{+}(k):=T_{A_{k}}$ and $T_{-}(k):=T_{-A_{k}}$. Notice that $T_{+}(k)$ admits rotational contact structures $\xi_{2n}$ for $n\geq 0$ and $T_{-}(k)$ admits rotational contact structures $\xi_{2n+1}$ for $n\geq 0$. There are additional contact structures $\eta_{n}$ on a parabolic torus bundle $T_{\pm}(k)$ for $n\in\mathbb{N}$.
@@ 1.7. Acknowledgements
We are grateful to Chris Wendl and Samuel Lisi for generous help in understanding their results and for crucial suggestions and input. We would also like to thank Mohammed Abouzaid, Inanc Baykur, Yakov Eliashberg, John Etnyre, Matt Hedden, Michael Hutchings, Olga Plamenevskaya, Laura Starkston, and Jeremy Van-Horn Morris for helpful conversations over the course of this project. The project was significantly helped by conferences which allowed us to collaborate in-person – we extend our thanks to the organizers of GSTGC 2022, Kylerec 2022, BIRS 2023 workshop on Interactions between Symplectic and Holomorphic Convexity in 4 Dimensions (23w5123), and the MSRI-PIMS 2022 Summer school on Floer Homotopy Theory. AR also acknowledges partial support from Georgia Tech and Louisiana State University, and the NSF grants DMS-1906414, DMS-2203312, and DMS-1907654. LW acknowledges support from NSF Grants DGE-2146752 and DMS-2303437 and IAS Giorgio and Elena Petronio Fellow II Fund.
@@ 2.1. Spinal open book decompositions and contact structures
    For every component of $\partial M_{P}$, there exist local coordinates $(\phi,t,\theta)\in S^{1}\times(-1,0]\times S^{1}$ on a collar neighborhood such that $\pi_{P}(\phi,t,\theta)=m\phi$ for some $m\in\mathbb{N}$. On every component of $\partial M_{\Sigma}$, there exist local coordinates $(s,\phi,\theta)\in(-1,0]\times S^{1}\times S^{1}$ such that $\pi_{\Sigma}(s,\phi,\theta)=(s,\phi)$. The number $m$ is called the multiplicity of $\pi_{P}$ at that boundary component of $M_{P}$. On the overlap between $M_{P}$ and $M_{\Sigma}$ the 2-torus coordinates $(\phi,\theta)$ agree, while the interval coordinates are related by $s=-t$.
@@
    and $\pi_{P}:M_{P}\to S^{1}$ is given by $[(\tau,p)]\to[\tau]$. Here, $\phi\in\operatorname{SMod}(P)$ is the monodromy in the spinal mapping class group of $P$. See §2.5 for the definition.

### v2

@@ Abstract.
Spinal open book decompositions provide a natural generalization of open book decompositions. We show that any minimal symplectic filling of a contact $3$-manifold supported by a planar spinal open book is deformation equivalent to the complement of a positive multisection in a bordered Lefschetz fibration, which generalizes a result of Wendl [Wen10c]. Along the way, we give an explicit local model for a non-compactly supported “singularity at infinity” in a generalized version of bordered Lefschetz fibrations, given by pseudoholomorphic foliations associated to the spinal open books. This provides new tools to classify symplectic fillings of a contact $3$-manifold that is not supported by an amenable spinal open book, by studying monodromy factorizations in the newly defined spinal mapping class group. As an application, we complete the classification of strong fillings of all parabolic torus bundles, and make progress towards classifying symplectic fillings of contact $3$-manifolds supported by non-planar open books.
@@ 1. Introduction
Finite-energy pseudoholomorphic foliations have played a central role in symplectic geometry for the past forty years. Using pseudoholomorphic foliations, Gromov [Gro85] and McDuff [McD90] initiated the study of the classification of symplectic $4$-manifolds. In contact geometry, early major applications include Eliashberg’s classification of symplectic fillings of the tight $3$-sphere [Eli91] and Hofer’s resolution of the Weinstein conjecture in several cases [Hof93]. These works inspired many developments later on in symplectic field theory (SFT), symplectic dynamics, and contact topology.
Our paper focuses on the problem of classification of symplectic fillings [^1] In this paper, we often omit strong in front of symplectic filling when it does not cause confusion. A strong symplectic filling $(X,\omega)$ of the contact manifold $(Y,\xi)$ is such that $\partial X=Y$ and $\omega|_{\text{nbd}\partial X}=d\lambda$ for some contact form $\lambda$ of $\xi$. of contact $3$-manifolds. For the tool of pseudoholomorphic foliations to apply, one has to restrict to the case where the leaves of the foliations are planar, due to index reasons. In the planar case, or more precisely, in the case of contact $3$-manifolds supported by planar open book decompositions, there are many past successes in the classification of their symplectic fillings, which is typically considered a very difficult problem, see [Wen10c, PVHM10a, Kal15, Pla12, Wan12, KL16] for examples. To go beyond the planar open book case, one key recent development is the technology of spinal open books initiated by Lisi, Van-Horn-Morris and Wendl in [LHMW18] and [LHMW20]. We discuss the current status and definition of spinal open books in §1.2 and §2.1. Our paper completes the topological understanding of a novel element in these studies: the role of “exotic fibers” in bordered Lefschetz fibrations that fill planar spinal open books. As a result, we are now able to give a complete description of symplectic fillings incorporating this new phenomenon. Our main theorems are stated in §1.3.
@@ 1.1. Scientific context
The main results of [Wen10c], [LHMW18] and [LHMW20] encapsulate a general theme of classifying minimal symplectic fillings of contact $3$-manifolds. The primary idea is to find planar pages in the contact $3$-manifold, which are genus-zero pages of its open book decomposition and which can be lifted to pseudoholomorphic curves in the symplectization. One then considers the moduli space of these pseudoholomorphic curves which turn out to foliate the symplectic fillings of the contact $3$-manifold, as in the main theorems of [Wen10c, LHMW18, LHMW20]. In [LHMW18] and [LHMW20], which generalize the results of [Wen10c], the authors show that in general, the moduli spaces contain two types of degenerations, namely singular fibers and exotic fibers. The exotic fibers do not appear in the setting of planar open books of [Wen10c], and the authors of [LHMW18] and [LHMW20] also only classify symplectic fillings in the restricted Lefschetz-amenable setting, where exotic fibers do not appear. They show that, in this situation, all minimal strong fillings are deformation equivalent to Stein fillings. Furthermore, the pseudoholomorphic foliation gives a bordered Lefschetz fibration structure on the strong filling, which induce the planar spinal open book on the boundary. These results already give vast generalization of the relative case of a result of McDuff [McD90] and unify the results of Gromov [Gro85], Eliashberg [Eli91], and McDuff [McD91], which classified the minimal symplectic fillings of the standard $3$-sphere and the standard lens spaces. In addition, the main result of [Wen10c] turns the classification problem for $3$-manifolds supported by planar open books into a monodromy factorization problem of the mapping classes supported in the interior of the page. This reduction to monodromy factorization was used extensively in [PVHM10a], [Kal15], [Pla12], [Wan12], and [KL16] to classify symplectic fillings of certain families of $3$-manifolds.
In this paper, we extend the results above to turn the problem of classifying symplectic fillings of a general planar spinal open book into a monodromy factorization problem, in a larger mapping class group that we call the spinal mapping class group of a surface. This mapping class group in particular contains diffeomorphisms of a surface that are not supported in the interior and will be defined in §2.5. We show that the presence of exotic fibers in minimal symplectic fillings forces the fillings to be complements of multisections in bordered Lefschetz fibrations. This phenomenon of exotic fiber can be considered as a “singularity at infinity” and has been largely unexplored. We expect our local model and description in terms of the spinal mapping class group to be also of interest for studying links of algebraic singularities in the spirit of [PS23, PVHM10b, KL16] and mirror symmetry.
@@ 1.2. Past structural theorem for planar spinal open books
Though the nomenclature and formal definition appeared later via the collaborations [LHMW18, LHMW20], the notion of spinal open books was first studied by Wendl in [Wen10c], where he used them to classify the symplectic and Stein fillings of the standard $3$-torus.
@@
[…]
@@
  • $)u\in\Pi^{-1}(\mathcal{M}_{\operatorname{\text{reg}}}(\widehat{J}$)) is an embedded $\widehat{J}-$holomorphic curve asymptotic to simply covered Reeb orbits;
  • $)u\in\Pi^{-1}(\mathcal{M}_{\operatorname{\text{sing}}}(\widehat{J}$)) is a nodal $\widehat{J}-$holomorphic curve asymptotic to simply covered Reeb orbits composing of two embedded curves intersecting transversely exactly once;
  • $)u\in\Pi^{-1}(\mathcal{M}_{\operatorname{\text{exot}}}(\widehat{J}$)) is an embedded $\widehat{J}-$holomorphic curve with one asymptotic end doubly covering a Reeb orbit and all other ends simply covering Reeb orbits.
@@
The above moduli spaces are referred to as the moduli space of regular, singular and exotic fibers, respectively. The pseudoholomorphic foliations given by Theorem 1.1 in the symplectic filling naturally give constraint to its boundary spinal open book. This leads to the seemingly technical definition of uniformity of a spinal open book, given in §2.1. As discussed in [LHMW20, Lemma 6.33], $\Pi|_{S_{i}}$ is an honest covering map if and only if $\mathcal{M}_{\operatorname{\text{exot}}}=\varnothing$. This condition on the spinal open book is called Lefschetz-amenable. In the Lefschetz-amenable setting, Lisi–Van-Horn-Morris–Wendl obtained the following result:
@@
The use of pseudoholomorphic foliations to provide rigidity results in symplectic geometry and dynamics is a well-established, powerful, and fruitful tool. The proof of Theorem 1.1 by Lisi–Van Horn-Morris-Wendl in [LHMW20] hinges upon the analysis of the SFT compactification of the moduli space of index two pseudoholomorphic foliations in a symplectic completion. Using this method to study finite-energy foliations for dynamical applications started with [Hof93] and many of the more general technical ingredients are developed in [HWZ96, HWZ95, BEH+03]. Recently, this strategy has been applied in [HSa11, HMSa15, CDR23] with great success for dynamical questions. For more geometric topological questions, this has been a long-term development of Wendl, starting primarily in [Wen10c, Wen13] by using finite-energy foliations to understand fillings of planar open books and develop filling obstructions. It is also the main technical tool in [DRGI16]. The results of [LHMW18, LHMW20] can be regarded as a continuation of the above mentioned projects. As previously mentioned, these ideas using foliations of pseudoholomorphic curves to classify symplectic $4$-manifolds date back to Gromov [Gro85] and McDuff [McD90]. In the usual game of using moduli spaces of pseudoholomorphic curves, there are mainly two concerns: compactness and transversality. The corresponding technical tools used for the setting of punctured finite-energy foliations are in [Wen10b] and [Wen10a], together with the intersection theory of punctured pseudoholomorphic curves developed in [Sie08, Sie11, Hut02], building upon [HWZ96, HWZ95, BEH+03] previously mentioned. We note that the automatic transversality criterion is satisfied for the moduli spaces of interest in the proof of Theorem 1.1 as a result of the curves considered being all of genus zero. To analyze the compactness problem, [LHMW20] carefully analyzes all possible configurations arising as the SFT limit and obtains the classification of curves as cited in Theorem 1.1.
The series of works [LHMW18, LHMW20] assembles multitudes of important techniques and geometric insights. However, due to its technical nature, the development of the spinal open book machinery, at the moment of their preprints, was ahead of its time and did not generate as many direct applications based on it. One exception is [BVHMLW15] finding contact $3$-manifolds with arbitrarily large Stein fillings. In contact and symplectic geometry of higher-dimensional manifolds, the technology of spinal open book has also inspired surprising applications. For example, Massot-Niederkrüger-Wendl found weakly but not Stein fillable manifolds [MNW13] using the spine removal surgery (see §2.7). Work of Bowden-Gironella-Moreno and Bowden-Gironella-Moreno-Zhou on the striking discovery of the abundance of tight non-fillable contact structures [BGM22, BGMZ22] also used the technology of spinal open books.
Via the full topological understanding of the associated pseudoholomorphic foliations with exotic fibers, we expect the technology of spinal open books and nearly Lefschetz fibrations to generate a new wave of activities. For example, nearly Lefschetz fibrations make a key contribution to the program of understanding fillings of links of sandwiched singularities by Plamenevskaya-Starkston [PS25]. In addition, in a sequel to the present article, the authors are working on classifying Stein and symplectic fillings of integer surgeries on various torus knots. Currently in the literature, a full classification of symplectic fillings for a given 3-manifold only exists for lens spaces [Lis08, CL23, ER21] and parabolic torus bundles; the latter done in this article combined with [LHMW20]. By applying our new technology, the sequel contributes yet another large class of full classification results previously inaccessible without the spinal open book tools.
@@ 1.3. Main results
We extend the results in the previous section to strong fillings of contact structures supported by uniform, non-Lefschetz-amenable spinal open books. The main contribution of our paper is to study the mysterious exotic fibers naturally arising in the non-Lefschetz-amenable situation, give local models and use them to classify fillings. As a result, we obtain a strong fillability criterion for non-Lefschetz-amenable planar spinal open books. This enlarges the types of singularities in the literature that is usually considered in the classification problem for compact symplectic manifolds in a significant way. See Remark 1.5 for other places where this type of singularity has been observed and relation to other works. In light of understanding these exotic fibers, we prove the following main result.
@@
We prove several technical results in §4 and §5 that allow us to conclude Theorem B. Firstly, Theorem 1.1 tells us that the completion of any strong filling of a partially planar spinal open book consists of a partition by regular, singular and exotic fibers. We establish a local model for exotic fibers in Proposition 4.9, which gives us a count of exotic fibers in the filling, as well as the monodromy around an exotic fiber. In terms of spinal open books, we state the following theorem.
@@ 1.4. Symplectic fillings of torus bundles
Ever since Eliashberg [Eli96] determined the strong fillability of contact structures on $T^{3}$, there has been extensive studies on the symplectic fillability of torus bundles. In particular, Stein and strong fillability of contact structures on torus bundles are completely determined except for negative hyperbolic ones, through extensive studies by [DG01, EH02, Gay06, VHM07, GL16, DL18, Chr21]. In addition, several classification results for these torus bundles are also known, see [EH02, Chr21, Kal15, Lis08, LHMW20, PVHM10a] for examples.
@@
where $\phi_{n}\colon\mathbb{R}$$\mathbb{Z}\to\mathbb{R}$ satisfies
@@
$\pm A_{k}=\pm\begin{pmatrix}1&k\\ 0&1\end{pmatrix}$
for some $k\in\mathbb{Z}$. We say a parabolic torus bundle $T_{A}$ is positive if $A$ is conjugate to $A_{k}$ and negative if $A$ is conjugate to $A_{k}$. Let $T_{+}(k):=T_{A_{k}}$ and $T_{-}(k):=T_{-A_{k}}$. Notice that $T_{+}(k)$ admits rotational contact structures $\xi_{2n}$ for $n\geq 0$ and $T_{-}(k)$ admits rotational contact structures $\xi_{2n+1}$ for $n\geq 0$. There are additional contact structures $\eta_{n}$ on a parabolic torus bundle $T_{\pm}(k)$ for $n\in\mathbb{N}$.
@@ 1.7. Acknowledgements
We are grateful to Chris Wendl and Samuel Lisi for generous help in understanding their results and for crucial suggestions and input. We would also like to thank Mohammed Abouzaid, Inanc Baykur, Georgios Dimitroglou Rizell, Yakov Eliashberg, John Etnyre, Matt Hedden, Helmut Hofer, Michael Hutchings, Olga Plamenevskaya, Laura Starkston, and Jeremy Van-Horn Morris for helpful conversations over the course of this project. The project was significantly helped by conferences which allowed us to collaborate in-person – we extend our thanks to the organizers of GSTGC 2022, Kylerec 2022, BIRS 2023 workshop on Interactions between Symplectic and Holomorphic Convexity in 4 Dimensions (23w5123), and the MSRI-PIMS 2022 Summer school on Floer Homotopy Theory. AR also acknowledges partial support from Georgia Tech and Louisiana State University, and the NSF grants DMS-1906414, DMS-2203312, and DMS-1907654. LW acknowledges support from NSF Grants DGE-2146752 and DMS-2303437 and IAS Giorgio and Elena Petronio Fellow II Fund.
@@ 2.1. Spinal open book decompositions and contact structures
    For every component of $\partial M_{P}$, there exist local coordinates $(\phi,t,\theta)\in S^{1}\times(-1,0]\times S^{1}$ on a collar neighborhood such that $\pi_{P}(\phi,t,\theta)=m\phi$ for some $m\in\mathbb{N}$. On every component of $\partial M_{\Sigma}$, there exist local coordinates $(s,\phi,\theta)\in(-1,0]\times S^{1}\times S^{1}$ such that $\pi_{\Sigma}(s,\phi,\theta)=(s,\phi)$. The number $m$ is called the multiplicity of $\pi_{P}$ at that boundary component of $M_{P}$. On the overlap between $M_{P}$ and $M_{\Sigma}$ the 2-torus coordinates $(\phi,\theta)$ agree, while the interval coordinates are related by s=ts=-t.
@@
    and $\pi_{P}:M_{P}\to S^{1}$ is given by $(\tau,p)]\to[\tau]$. Here, $\phi\in\operatorname{SMod}(P)$ is the monodromy in the spinal mapping class group of $P$. See §2.5 for the definition.